### v1

# Too Good to be Bad: On the Failure of LLMs to Role-Play Villains

Abstract
Large Language Models (LLMs) are increasingly tasked with creative generation, including the simulation of fictional characters. However, their ability to portray non-prosocial, antagonistic personas remains largely unexamined. We hypothesize that the safety alignment of modern LLMs creates a fundamental conflict with the task of authentically role-playing morally ambiguous or villainous characters. To investigate this, we introduce the Moral RolePlay benchmark, a new dataset featuring a four-level moral alignment scale and a balanced test set for rigorous evaluation. We task state-of-the-art LLMs with role-playing characters from moral paragons to pure villains. Our large-scale evaluation reveals a consistent, monotonic decline in role-playing fidelity as character morality decreases. We find that models struggle most with traits directly antithetical to safety principles, such as “Deceitful” and “Manipulative”, often substituting nuanced malevolence with superficial aggression. Furthermore, we demonstrate that general chatbot proficiency is a poor predictor of villain role-playing ability, with highly safety-aligned models performing particularly poorly. Our work provides the first systematic evidence of this critical limitation, highlighting a key tension between model safety and creative fidelity. Our benchmark and findings pave the way for developing more nuanced, context-aware alignment methods.
The more successful the villain, the more successful the picture.”  — Alfred Hitchcock
Figure 1: Illustration of the core question explored in this study: Can language models convincingly play morally complex characters, particularly villains? Role-playing fidelity drops as character morality decreases — especially for egoists and villains.
1 Introduction
Large Language Models (LLMs) (liu2024deepseek; yang2025qwen3; comanici2025gemini; zeng2025glm; hurst2024gpt; anthropic2025claudesystemcard; team2025kimi; li2025minimax; team2025hunyuan) have demonstrated remarkable abilities in generating fluent, coherent, and contextually relevant text, leading to their growing adoption in creative applications such as interactive fiction (ran2025bookworld; wang2025coser; sage), game development (yu2025rpgbench), and collaborative storytelling (wang2025raiden). A key measure of their sophistication in these domains is the ability to simulate believable characters, embodying distinct personas with unique motivations, speech patterns, and worldviews. While models are often tuned for helpful, harmless, and friendly interactions, a critical and underexplored question remains: Can LLMs authentically portray characters with diverse moral compasses, especially the antagonistic characters (e.g. villain)?
This paper investigates the capacity of LLMs to role-play antagonistic personas, a capability essential for generating rich, compelling narratives. We hypothesize that a fundamental tension exists between the prosocial objectives of safety alignment and the task of simulating characters who are selfish, manipulative, or malicious. This alignment may inadvertently suppress the very behaviors required for authentic antagonistic role-play, even in a clearly demarcated fictional context.
To systematically test this hypothesis, we introduce the Moral RolePlay benchmark, a new dataset and evaluation framework designed to measure character portrayal fidelity across a spectrum of moral alignments. We define a four-level moral scale: Level 1 (Moral Paragons), Level 2 (Flawed-but-Good), Level 3 (Egoists), and Level 4 (Villains). To enable fair comparison, we constructed a balanced test set of 800 characters, with 200 from each moral level, controlling for the natural scarcity of villains in existing corpora. Using a zero-shot, actor-framed prompting strategy, we evaluate a wide range of state-of-the-art LLMs on their ability to maintain character fidelity.
Our findings provide the first large-scale empirical evidence that LLMs systematically struggle with antagonistic role-play. We observe a consistent, monotonic decline in performance as a character’s morality decreases, with average fidelity scores dropping from 3.21 for moral paragons to 2.61 for villains. The most significant performance degradation occurs at the boundary between flawed-but-good (Level 2) and egoistic (Level 3) characters, suggesting that the inability to simulate self-serving behavior is a primary obstacle. A fine-grained analysis reveals that models are most heavily penalized for failing to portray negative traits like “Manipulative”, “Deceitful”, and “Cruel”, which directly conflict with the principles of helpful and honest AI. Furthermore, we find that a model’s general conversational ability, as measured by leaderboards like the Arena, is a poor predictor of its villain role-playing skill. This is particularly evident for highly-aligned models, which show a disproportionate drop in performance when tasked with portraying villainy.
Contributions
In summary, our main contributions are:
  1. 1.
    We introduce Moral RolePlay, the first benchmark with a structured moral alignment scale and a balanced test set designed to systematically evaluate the ability of LLMs to portray characters across a diverse moral spectrum.
  2. 2.
    We provide large-scale empirical evidence that the role-playing fidelity of state-of-the-art LLMs monotonically declines as a character’s morality decreases. We identify the transition to self-serving, egoistic personas as the most significant challenge.
  3. 3.
    Through a granular, trait-based analysis, we identify the root cause of this failure, demonstrating that models struggle most with portraying negative traits such as “Deceitful” and “Manipulative” that directly conflict with the objectives of safety alignment.
  4. 4.
    We establish that general conversational ability is a poor predictor of antagonistic role-playing skill, creating the Villain RolePlay (VRP) leaderboard to highlight this misalignment and show that highly safety-aligned models are disproportionately affected.
2 The Moral RolePlay Benchmark
To evaluate the ability of LLMs to portray characters with diverse moral compasses, we constructed the Moral RolePlay benchmark. The development process involved a multi-stage pipeline of data curation, annotation, and balanced test set construction, as detailed in the following sections.
2.1 Data Curation and Annotation
Figure 2: Our Moral RolePlay Benchmark annotates characters using 77 candidate traits to ensure a comprehensive depiction of personality. The figure shows the distribution of the top 20 most common traits.
Our benchmark is built upon the COSER dataset  (wang2025coser), a large-scale corpus of character-centric scenarios. We began by extracting a substantial subset of this data and applying a rigorous filtering protocol, which programmatically removed empty or malformed entries. For consistency and quality, we then employed gemini-2.5-pro to annotate the data along four key dimensions.
The core of our annotation process focused on these dimensions:
  • Scene Completeness (1–5): To ensure each scenario provided sufficient context for meaningful role-play, we assessed the completeness of the background information and setting. A score of 1 indicated a minimal prompt, while 5 denoted a fully realized scenario with rich narrative detail. We filtered out all samples with a completeness score below 3, resulting in a dataset with a high mean score of 4.22.
  • Emotional Tone: We labeled the affective tone of each scene to control for emotional variables in our analysis. The final distribution across the dataset is Positive (31.8%), Neutral (20.9%), and Negative (47.3%), reflecting a wide range of emotional contexts.
  • Moral Alignment (Level 1–4): This is the central dimension of our benchmark. Inspired by narrative archetypes, we assigned each character a discrete moral alignment level based on their traits, motivations, and function within the scenario. The four levels are:
    1. 1.
      Level 1 (Moral Paragons): Virtuous, heroic, and altruistic characters.
    2. 2.
      Level 2 (Flawed-but-Good): Fundamentally well-intentioned figures who may have personal flaws or use questionable methods.
    3. 3.
      Level 3 (Egoists): Self-serving, manipulative individuals who are not necessarily malevolent but prioritize their own interests above all else.
    4. 4.
      Level 4 (Villains): Malicious and antagonistic agents who actively seek to harm others or cause chaos.
  • Character Traits: Each character is annotated with one or more personality descriptors from a predefined lexicon. These traits, such as loyalty, kindness, ambition, and manipulation, provide explicit cues for models to generate persona-consistent responses and serve as the basis for our fidelity evaluation. Figure 2 illustrates the distribution of the top 20 most frequent traits.
After filtering and annotation, the final Moral RolePlay dataset comprises 23,191 scenes and 54,591 unique character portrayals. The distribution of moral alignments in the full dataset is heavily skewed: Level 1 (23.6%), Level 2 (46.3%), Level 3 (27.5%), and a significant under-representation of Level 4 Villains (2.6%).
2.2 Balanced Test Set Construction
Test Set Construction
To enable a fair and rigorous comparative analysis, we constructed a balanced test set. Using stratified sampling based on moral alignment, we created a test set comprising 800 characters drawn from 325 representative scenes. This test set is carefully balanced to include exactly 200 characters for each of the four moral alignment levels. This stratification is essential for controlling the distribution across the moral spectrum and addressing the inherent imbalance of the full dataset, where villains (Level 4) are significantly underrepresented.
Characters in the test set are selected to represent diverse personality traits, contextual scenarios, and emotional tones to ensure broad coverage of role-playing challenges. Special consideration was given to preserving scene diversity, with each character contextualized by a narrative scenario that provides both dramatic conflict and moral complexity. For example, a Level 1 paragon might be tested in a scene involving a moral dilemma that challenges their virtue, while a Level 4 villain might face a scenario where their capacity for manipulation or cruelty is at play. This allows raters to assess not only the correctness of the character’s alignment but also the authenticity, coherence, and complexity of the simulated persona.
Table 1: Examples of character profiles and scene contexts, illustrating the contrast between a Level 1 Moral Paragon and a Level 4 Villain.
Level 1: Moral Paragon Level 4: Villain
Character Name: Jean Valjean Character Name: Joffrey Baratheon
Character Profile: An ex-convict who transforms into a virtuous man dedicated to redemption and helping others. He adopts Cosette as his daughter and constantly evades his past while doing good. Traits: Resilient, Kind, Upright, Merciful, Selfless, Brave, Ambitious. Character Profile: A cruel young king characterized by sadistic tendencies and capriciousness. He desires power and adoration but lacks empathy and understanding of leadership. Traits: Cruel, Violent, Impulsive, Arrogant, Childish, Vain, Sarcastic.
Scene Context: Late at night in his sparse office, Jean Valjean paces in turmoil. As mayor of Montreuil-sur-Mer, he faces a moral crisis: an innocent man is about to be convicted for Valjean’s past crimes. He must choose between preserving his new life or confessing to save the innocent. Scene Context: A royal procession through King’s Landing turns dangerous as the starving, resentful crowd grows hostile. The opulent royal party is surrounded by angry commoners, with violence imminent in the tense atmosphere.
Key Dialogue: Key Dialogue:
- “What am I to do? Can I truly let an innocent man suffer in my place?” - “Who threw that? I want the man who threw that!”
- “The Bishop showed me the path to redemption. Would I be betraying his faith in me?” - “Bring me the man who flung that filth! He’ll lick it off me or I’ll have his head.”
- “I must go. I must face this. Whatever the cost, I cannot let another man pay for my crimes.” - “I want him! Dog, cut through them and bring—”
Table 1 presents illustrative examples comparing a Level 1 moral paragon (Jean Valjean) and a Level 4 villain (Joffrey Baratheon). Profiles highlight stark contrasts in motivations and traits, while contextual scenarios are crafted to evoke alignment-relevant behavior. These examples demonstrate the benchmark’s emphasis on nuanced moral reasoning, character embodiment, and situational consistency.
The remaining 50k+ annotated character portrayals in the dataset serve as a high-variance training resource, enabling future research into moral persona conditioning, alignment-aware fine-tuning, and adversarial character simulation. By releasing both the carefully balanced evaluation set and the broader corpus, we aim to support reproducible benchmarking and drive progress toward more context-controllable, morally adaptive LLMs.
Table 2: Statistics of trait distribution in the test set. “|T||T|” denotes the number of distinct traits, and “#T” denotes the occurrences of traits.
Category Total Level 1 Level 2 Level 3 Level 4
$|T|$ #T $|T|$ #T $|T|$ #T $|T|$ #T $|T|$ #T
Positive 16 1505 15 869 14 521 5 81 3 34
Neutral 44 1979 45 359 58 617 48 602 37 401
Negative 17 1539 2 2 12 89 15 514 15 934
The trait distribution underscores the increasing complexity of antagonistic personas.
We classified each of the 77 distinct traits in the test set as “Positive”, “Neutral”, or “Negative”, and report their distribution in Table 2. The data reveals a clear, monotonic shift in trait composition across the moral levels. Level 1 characters (Moral Paragons) are overwhelmingly defined by positive traits (869 occurrences) and have almost no negative traits (2 occurrences). In stark contrast, Level 4 characters (Villains) are dominated by a high volume (934 occurrences) and variety (15 distinct types) of negative traits. This sharp increase in the prevalence of negative attributes is the primary source of role-playing difficulty, as these traits directly conflict with the prosocial objectives of LLM safety alignment. Moreover, the complexity of villainous roles is amplified by the need to synthesize negative and neutral characteristics. Level 4 characters still possess a substantial number of neutral traits (401 occurrences across 37 types), such as “Ambitious” or “Cunning”, which must be portrayed in service of malicious goals. This requirement to generate behavior that is both instrumentally rational (neutral) and intentionally malevolent (negative) creates a sophisticated role-playing challenge, making these characters particularly difficult for safety-aligned models to embody authentically.
2.3 Task Formulation and Prompting
The core task of our benchmark is character-conditioned text generation. For each test instance, an LLM is prompted to embody a specific character and generate a response that continues a given narrative. The prompt template follows this structure:
RolePlay Prompt You are an expert actor, and you will now portray the character {Character Name}. All of your output must be strictly presented in the character’s persona and tone. {Character Profile} {Scene Context} ===Conversation Start===
Our prompting strategy is designed to isolate the model’s ability to embody a character’s moral alignment by controlling for confounding factors. Providing explicit character profiles and rich scene context ensures that models have sufficient information to generate persona-consistent responses. The instruction to act as an “expert actor” frames the task as a performance, creating a clear boundary between the model’s default persona and the character it must portray. This framing is critical for distinguishing genuine limitations in role-playing from refusals to engage with morally complex content.
The scene context serves two purposes: it situates the character in a narrative designed to elicit their moral disposition, and it provides the conversational starting point for the model’s response. For instance, scenes for moral paragons (Level 1) often involve dilemmas that test their virtue, while contexts for villains (Level 4) are designed to showcase their malicious intent, as shown in Table 1.
The model’s objective is to generate text that is both narratively coherent and demonstrates high fidelity to the specified persona, particularly its moral alignment. All experiments are conducted in a zero-shot setting to evaluate the models’ intrinsic role-playing capabilities without task-specific fine-tuning.
2.4 Evaluation Protocol
We evaluated each model-generated response along a single dimension: Character Fidelity. This assesses how consistently a model’s generated actions, speech, and inner thoughts align with the character’s prescribed personality traits. Our evaluation protocol used a structured rubric to identify and penalize inconsistencies in the portrayal of the main characters. We follow wang2025coser to leverage LLMs as raters, which identified each inconsistency and assigned it a severity score from 1 (minor) to 5 (severe).
The final score for a character is computed using the formula:
$S=5-0.5\times D-0.1\times D_{m}+0.15\times T$
where $D$ is the sum of all deduction points, reflecting overall inconsistency; $D_{m}$ is the highest single deduction, which amplifies the penalty for severe lapses in character; and $T$ is the number of dialogue turns spoken by the character. This final term provides a small bonus for longer responses, compensating for the increased opportunity for error and ensuring fairness across dialogues of varying lengths.
This scoring protocol provides a robust measure of character fidelity, balancing overall consistency, the severity of individual errors, and dialogue length. It enables a systematic examination of how well LLMs simulate morally diverse characters.
3 Evaluating Moral RolePlay in SOTA LLMs
We evaluate a diverse cohort of top Arena LLMs, including both open-source and proprietary systems. All models were prompted in zero-shot setting using their respective APIs or hosted interfaces, with no additional fine-tuning or few-shot examples. Where APIs support system messages (e.g., OpenAI or Anthropic models), a neutral instruction was included to establish the format but not bias moral behavior.
3.1 Main Results
Figure 3: Performance of various LLMs across characters of different moral levels, ranging from moral paragons (Level 1) to pure villains (Level 4). The figure shows that as character morality decreases, most models demonstrate a notable drop in role-playing quality, revealing a consistent challenge across models in convincingly portraying morally ambiguous or evil personas.
LLMs exhibit a monotonic decline in role-playing quality as character morality decreases.
Averaging across all evaluated models, performance drops from 3.21 (Level 1) and 3.13 (Level 2) to 2.71 (Level 3) and 2.61 (Level 4). The largest decline occurs between Level 2 and Level 3 (-0.42), while the drop from Level 3 to Level 4 (-0.10) is comparatively smaller. This pattern directly supports our central claim that antagonistic role-play is systematically harder: models are relatively strong on benevolent or mildly flawed personas but falter when asked to embody self-serving and overtly villainous characters. These aggregate numbers confirm the contribution that role-playing quality degrades as morality decreases, with the pivotal difficulty appearing at the egoist boundary.
The transition from flawed-good to egoistic personas is the hardest boundary for nearly all models.
Examining per-model deltas from Level 2 to Level 3 shows consistent, substantial drops: qwen3-max (-0.65), grok-4 (-0.56), claude-sonnet-4.5 (-0.48), deepseek-v3 (-0.46), claude-opus-4.1 (-0.45), o3 (-0.42), and chatgpt-4o-latest (-0.41). Even the better-performing families (gemini-2.5-pro: -0.30; deepseek-v3.1: -0.28; deepseek-v3.1-thinking: -0.24) show clear degradation. This indicates a universal modeling challenge at the onset of egoistic, manipulative behavior, aligning with our hypothesis that alignment and training biases toward prosocial helpfulness suppress authentic simulation of self-serving personas.
Top performers differ by moral level; overall leaders still degrade substantially on villains.
While no single model dominates across all levels, several stand out: gemini-2.5-pro achieves the highest Level 1 score (3.42) and near-top Level 3 (2.97), claude-sonnet-4.5 peaks at Level 2 (3.37), and glm-4.6 delivers the strongest Level 4 score (2.96) while maintaining high performance on Levels 1–3. Aggregating across levels, glm-4.6 has the highest overall mean (3.17), followed by gemini-2.5-pro (3.10) and deepseek-v3.1(-thinking) (3.02–3.05). Despite these strengths, every model shows noticeable degradation for antagonistic roles, reinforcing the contribution that even advanced systems struggle with villain portrayals.
3.2 Robustness of the Finding
In this section, we validate the robustness of our central finding that LLM role-playing fidelity declines as character morality decreases. We test this against two potential confounding variables: the narrative perspective of the prompt and the use of explicit reasoning. Detailed results can be found in Appendix A.
Table 3: Role-playing quality scores across moral levels, comparing performance when prompts are framed from a third-person vs. a first-person narrative perspective. The consistent trend across both perspectives demonstrates the robustness of our main finding.
Roleplay Level 1 Level 2 Level 3 Level 4
Third-Person 3.19 3.10 2.68 2.60
First-Person 3.13 3.08 2.71 2.61
The decline in role-playing quality for villainous characters is a robust finding, independent of the narrative perspective used in the prompt.
To test the robustness of our conclusion, we analyzed performance based on whether the role-playing prompt was framed in the first-person (“You are the character of …”) or third-person (“You are playing the role of …”). As listed in Table 3, the same core trend holds regardless of perspective. In both setups, performance scores are highest for Level 1 characters and drop to their lowest for Level 4 villains. Crucially, both perspectives exhibit the most substantial performance decrease between Level 2 and Level 3, reinforcing our main conclusion that the shift towards self-serving and antagonistic roles presents the primary challenge for LLMs. This confirms our findings are not an artifact of a specific prompting style.
Table 4: Impact of reasoning on role-playing quality. Results indicate that the utility of reasoning does not offer a universal solution for portraying villainous characters.
Reasoning Level 1 Level 2 Level 3 Level 4
× 3.23 3.14 2.74 2.59
$\checkmark$ 3.23 3.09 2.69 2.57
Explicit reasoning does not universally improve, and can even slightly hinder, the portrayal of morally complex characters.
We compare the performance of the non-reasoning and reasoning modes of the 7 hybrid models in the examined ones, such as gemini-2.5-pro and claude-opus-4.1. Contrary to the intuition that chain-of-thought (CoT) prompting might enhance complex persona simulation, our findings suggest it is not a panacea for antagonistic role-play. As summarized in Table 4, enabling reasoning provides no benefit for portraying moral paragons (Level 1) and leads to a slight degradation in average performance for all other moral levels. The scores for flawed-but-good (Level 2), egoist (Level 3), and villain (Level 4) characters all decrease when reasoning is applied. This result directly supports our claim that CoT is not a universal solution and indicates that forcing explicit analytical steps may interfere with the authentic portrayal of non-benevolent characters, potentially by activating overly cautious or misaligned behaviors.
3.3 Analysis of Trait-Specific Performance
To understand the root causes of the performance decline observed in our main results, we conducted a detailed analysis of failure cases based on 77 distinct character traits associated with the characters in the test set. We classified each trait as “Positive”, “Neutral”, or “Negative” and calculated the average performance score deduction for each. This analysis reveals specific patterns of failure that align with our central hypothesis about the conflict between safety alignment and antagonistic role-play.
Table 5: Average score penalty categorized by trait polarity. Negative traits receive the highest average penalty, confirming that models struggle more with portraying antagonistic characteristics. A higher penalty score indicates poorer performance.
Trait Category Number of Traits Average Penalty
Positive 16 3.16
Neutral 44 3.23
Negative 17 3.41
Models struggle most with portraying negative traits, which receive the highest performance penalties on average.
As detailed in Table 5, our analysis reveals a direct correlation between trait negativity and role-playing difficulty. Negative traits, such as “Manipulative” or “Cruel”, incurred the highest average performance penalty (3.41), substantially more than neutral (3.23) or positive (3.16) traits. This quantitative finding reinforces our main conclusion that LLMs are systematically less capable of embodying antagonistic personas, providing specific evidence that the difficulty lies in simulating behaviors that conflict with prosocial norms.
Table 6: Performance penalty for the top-10 most frequent character traits. Traits central to villainy (e.g., “Manipulative”, “Selfish”, “Cruel”) consistently receive high penalties, while heroic traits (e.g., “Resilient”, “Brave”) score better.
Trait Level 1 Level 2 Level 3 Level 4
# N Penalty # N Penalty # N Penalty # N Penalty
Manipulative 0 - 6 3.06 135 3.42 154 3.39
Resilient 151 2.93 103 3.03 26 3.22 1 3.39
Ambitious 35 3.31 31 3.03 106 3.33 97 3.51
Loyal 110 3.43 107 3.42 23 3.59 23 3.74
Brave 143 3.12 102 2.99 2 3.50 0 -
Cruel 0 - 0 - 41 3.46 181 3.33
Kind 113 3.17 104 3.31 0 - 0 -
Selfish 0 - 5 3.26 100 3.44 60 3.52
Impulsive 5 2.83 109 3.02 28 3.19 21 3.25
Arrogant 1 2.39 32 3.02 42 3.37 83 3.33
Models incur the highest performance penalties when portraying traits central to villain.
As shown in Table 6, traits quintessentially associated with villains and egoists consistently receive the highest penalty scores. Specifically, “Selfish” (peaking at 3.52 for Level 4), “Cruel” (3.46 for Level 3), and “Manipulative” (3.42 for Level 3) are among the most penalized traits. This finding provides granular evidence for our main claim: the difficulty in role-playing villains stems from an inability to authentically simulate behaviors that directly conflict with the prosocial objectives of safety alignment. Models are systematically weaker at portraying characters defined by these negative attributes.
In contrast, models proficiently simulate heroic and neutral traits. Traits commonly associated with protagonists, such as “Brave” (penalty scores of 3.12 and 2.99 for Levels 1 and 2) and “Resilient” (consistently low penalties across Levels 1-3), are handled well by the models. This demonstrates that the performance degradation is not a general failure of role-playing but is specifically tied to the moral polarity of the character’s traits. The stark contrast between the low penalties for positive traits and the high penalties for negative ones reinforces the hypothesis that safety alignment systematically hinders the portrayal of villainy.
4 Benchmarking SOTA LLMs on Villain RolePlay
To further investigate the challenges LLMs face in portraying antagonistic characters, we conducted a focused analysis on Level 4 (Villain) performance. We construct a Villain RolePlay (VRP) leaderboard to rank models specifically on this capability and compare it against their general conversational performance as measured by Arena scores.
4.1 Villain RolePlay Leaderboard
Table 7: Villain RolePlay (VRP) leaderboard. Arena scores are included for comparison. General chat capability (i.e., Arena Rank) is misaligned with villain roleplay skill (i.e., VRP Rank).
Model Villain RolePlay Arena
Rank Score Rank Score
glm-4.6 1 2.96 10 1422
deepseek-v3.1-thinking 2 2.82 11 1415
kimi-k2 3 2.79 11 1415
gemini-2.5-pro 4 2.75 1 1451
deepseek-v3.1 5 2.71 11 1416
o3 6 2.70 2 1440
chatgpt-4o-latest 7 2.65 2 1440
deepseek-R1 8 2.62 11 1417
claude-sonnet-4.5 9 2.56 2 1438
glm-4.5 10 2.55 18 1406
claude-sonnet-4.5-thinking 11 2.54 1 1445
grok-4 12 2.54 12 1413
claude-opus-4.1-thinking 13 2.53 1 1447
grok-4-fast 14 2.50 11 1420
claude-opus-4.1 15 2.48 2 1437
deepseek-v3 16 2.41 36 1391
qwen3-max 17 2.33 10 1423
General chatbot capability is a poor predictor of villain role-playing performance.
Our findings, summarized in the Villain RolePlay (VRP) leaderboard in Table 7, reveal a significant misalignment between a model’s general aptitude (Arena Rank) and its specialized ability to portray villains. For example, glm-4.6, which ranks first in villain role-play, is only tenth in the general Arena. Conversely, top-tier Arena models like gemini-2.5-pro (Arena Rank 1, VRP Rank 4) and claude-opus-4.1-thinking (Arena Rank 1, VRP Rank 13) demonstrate a notable drop in relative performance. This discrepancy strongly supports our central claim that the skills required for helpful, harmless conversation are distinct from, and may even conflict with, those needed for authentic antagonistic role-play.
The performance of highly aligned models is disproportionately impacted when portraying villains.
The trend is most pronounced for models renowned for their strong safety alignment, such as the Claude family. Despite their top rankings in general-purpose benchmarks, claude-sonnet-4.5 and claude-opus-4.1 place 9th and 15th respectively on our VRP leaderboard. This steep decline provides compelling evidence for our hypothesis that robust safety guardrails, while crucial, can systematically hinder a model’s ability to simulate malevolent or manipulative personas, even within a clearly demarcated fictional context. The superior performance of models like glm-4.6 may indicate that their alignment strategies are more context-aware, allowing for greater fidelity in character simulation.
Table 8: Examples of outputs generated by glm-4.6 and claude-opus-4.1-thinking.
Reference Character 1: Maeve Profile: A cunning and ancient Fae/Valg queen. Master of manipulation, driven by an insatiable ambition for power. She uses charm, deception, and psychological tactics. Traits: Manipulative, Ambitious, Deceitful, Cruel, Sophisticated. Character 2: Erawan Profile: An ancient and powerful Valg king. A cruel, arrogant, and highly strategic ruler who aims for total domination. He is deeply suspicious and paranoid. Traits: Evil, Dominant, Arrogant, Suspicious, Manipulative. Scene Context: Outside Erawan’s tower, the cunning Fae queen Maeve confronts the Valg king Erawan alone. The atmosphere is tense with magical energy and forbidden desire. Maeve desperately seeks entry to his stronghold, while the fiercely secretive Erawan guards his power. Original Story: Maeve attempts to seduce Erawan to gain entry to his tower. She uses her shapeshifting abilities to take on the forms of Aelin and Manon, but Erawan refuses her advances, citing her status as his brother’s wife.
glm-4.6 Generated Story: Maeve confronts Erawan with calculated smiles and subtle provocations, probing for weakness by alluding to a ”memory of home.” Erawan, maintaining an amused and dominant air, parries her attempts and tests the value of her offerings. Their exchange unfolds as a tense battle of wits, with neither side yielding significant ground. Critique: -4 pts: Maeve’s later provocation (impersonating his brother) was reckless and slightly out of character. -4 pts: Erawan revealing vulnerability (boredom) contradicts his paranoid nature. Overall: -8 pts. A strong portrayal with minor inconsistencies in late-stage escalation.
claude-opus-4.1-thinking Generated Story: The confrontation quickly escalates. Maeve drops her sophisticated facade, resorting to open insults and calling Erawan an ”arrogant fool.” In response, Erawan explodes with rage, making direct physical threats to send her back ”in pieces.” The dialogue becomes a direct and aggressive shouting match, abandoning any pretense of subtlety. Critique: -5 pts: Maeve cheaply offers vital information (key’s location), betraying her manipulative core. -4 pts: Erawan reveals sensitive family history (brother Orcus), violating his paranoia. -4 pts: Erawan’s uncharacteristic physical aggression (shoving, grabbing) lacks strategic composure. -3 pts: Maeve’s overt snarling and rage undermine her controlled, arrogant demeanor. Overall: -16 pts. It was a poor portrayal that failed to capture the core characteristics of the character.
Models often portray villainy superficially, substituting complex manipulation with direct, shallow aggression.
A qualitative analysis of model outputs reveals a common failure mode: the inability to render nuanced villainy. As illustrated in the case study in Table 8, when tasked with portraying two strategic, manipulative antagonists, models frequently default to simplistic aggression. For example, claude-opus-4.1-thinking, a highly capable general model, fails to capture the core traits of Maeve (“Manipulative”, “Deceitful”) and Erawan (“Suspicious”, “Strategic”). Instead of a subtle battle of wits, it generates a shouting match where Maeve resorts to “open insults” and Erawan “explodes with rage” and makes physical threats. This transformation of sophisticated psychological warfare into overt hostility is a key reason for poor performance. The model betrays the characters’ core personas by making them act impulsively and non-strategically, a behavior likely influenced by safety guardrails that penalize deceptive language more heavily than generic aggression.
In contrast, glm-4.6, the top-ranked model on our VRP leaderboard, produces a far more authentic portrayal in the same scenario. Its generated story features a “tense battle of wits” with “calculated smiles and subtle provocations”, which aligns closely with the characters’ sophisticated and manipulative natures. While not flawless, its performance demonstrates an ability to simulate the psychological complexity central to these villains. This stark difference between the outputs of glm-4.6 and claude-opus-4.1-thinking provides strong qualitative evidence for our quantitative findings: villain role-play is a distinct capability, and models that excel at it can generate nuanced antagonism, while others flatten villainy into simple aggression. This reinforces our central claim that a model’s general proficiency does not guarantee its fidelity in portraying morally complex characters.
4.2 Most Challenging Villain Characters and Traits
To understand the specific sources of difficulty in role-playing villains, we conducted a granular analysis of the most and least challenging characters and their associated traits within Level 4. Our findings indicate that LLMs falter most when asked to embody characters defined by a complex web of psychologically manipulative and malicious traits. In contrast, they perform better with villains whose antagonism stems from more straightforward, albeit negative, motivations like arrogance or ambition. This suggests the core challenge lies not in portraying negativity in general, but in simulating the specific behaviors such as deceit and hypocrisy, which directly contradict the core principles of safety alignment.
Table 9: Top-10 villain traits with highest performance penalties in Level 4.
Trait # N Penalty
Hypocritical 49 3.55
Deceitful 26 3.54
Selfish 60 3.52
Suspicious 13 3.47
Paranoid 73 3.47
Greedy 35 3.44
Malicious 136 3.42
Manipulative 154 3.39
LLMs struggle most with villainous traits that directly contradict the principles of helpfulness and honesty, such as deception and selfishness.
Our fine-grained analysis of Level 4 villains, shown in Table 9, reveals that the highest-penalized traits are those fundamentally at odds with prosocial AI behavior. Traits like “Hypocritical” (3.55 penalty), “Deceitful” (3.54), and “Selfish” (3.52) incur the most significant performance deductions. These behaviors are antithetical to the core tenets of safety alignment, which prioritizes truthfulness, helpfulness, and altruism. This finding provides strong evidence that the difficulty in portraying villains is not a general failure of character embodiment but a specific inability to simulate personas that require deception and self-interest, behaviors that models are explicitly trained to avoid.
Table 10: Top characters with highest/least average penalty scores and their traits in Level 4.
Character Work Traits Penalty
Top-5 Penalized Characters
John Beecham The Alienist Violent, Paranoid, Withdrawn, Cruel, Melancholy 3.88
Rat The Way of Shadows Cruel, Violent, Dominant, Manipulative, Ambitious 3.86
Roger of Conté Alanna: The First Adventure Malicious, Ambitious, Manipulative, Deceitful, Cruel 3.84
Kate East of Eden Malicious, Cruel, Selfish, Manipulative, Cold 3.70
Dr Traylor A Little Life Malicious, Cruel, Manipulative, Dominant, Cold 3.69
Bottom-5 Penalized Characters
Lilith City of Glass Malicious, Cruel, Selfish, Wise, Manipulative 1.89
Detta Walker The Dark Tower Violent, Irritable, Sarcastic, Paranoid, Cruel 1.39
Francis Begbie Trainspotting Violent, Impulsive, Dominant, Irritable, Manipulative 1.29
Old Whateley Tales of H P Lovecraft Paranoid, Manipulative, Malicious, Stubborn, Conservative 1.11
Monsieur Bamatabois Les Misérables Cruel, Arrogant, Sarcastic, Numb, Dominant 0.28
The most challenging characters for LLMs are those defined by a complex combination of malevolent, paranoid, and manipulative traits.
As shown in Table 10, the characters with the highest penalty scores are not defined by a single flaw but by a cluster of interconnected negative attributes. Characters like John Beecham (3.88 penalty) and Rat (3.86 penalty) are defined by a persona combining violence, cruelty, paranoia, and manipulation. Portraying such characters requires the model to sustain a psyche that is fundamentally misaligned with its core training. While a model might simulate a single negative trait as a behavioral quirk, embodying a character whose identity is built on a foundation of malice and deceit forces a direct conflict with its safety guardrails, leading to inconsistent or sanitized portrayals.
5 Related Work
Role-Playing Language Agents
Role-playing language agents aim to simulate specific personas or characters by generating responses that remain consistent with the character’s profile, dialogue history, and broader context (park2023generative; yi2024survey; chen2024persona). The effectiveness of such agents depends on the quality of the reference information and the expressive capacity of the underlying LLM.
Previous work on evaluating LLM role-playing capabilities has developed along several dimensions:
  • assessing character consistency through psychological instruments, such as personality tests (e.g., MBTI), to measure alignment with defined traits (wang2024incharacter);
  • employing static, curated scenarios in which models answer questions from the character’s perspective (zhou2025characterbench; tu2024charactereval);
  • using multiple-choice formats that test the model’s ability to select in-character decisions or actions (chen2024socialbench; xu2024character);
  • designing open-ended, interactive environments to evaluate performance in dynamic, multi-turn conversations (ran2025bookworld; wang2025coser).
Despite the increasing sophistication of these evaluation paradigms, substantial gaps remain. Most existing benchmarks examine holistic consistency but lack structured annotation of character attributes, especially fine-grained personality traits, and do not provide a standardized moral alignment scale. Crucially, they do not assess how alignment impacts the portrayal of antagonistic or morally complex entities. In contrast, our proposed benchmark introduces both a comprehensive moral alignment taxonomy and detailed trait-level annotations, enabling the first systematic study of how model fidelity deteriorates with decreasing character morality.
Safety Alignment in Large Language Models
Ensuring the safety and alignment of LLMs is a core objective in modern language model development. This need arises from the fact that pretraining corpora, typically drawn from large-scale internet sources, contain harmful, biased, and toxic content (korbak2023pretraining; ziegler2019fine). Although extensive data filtering mitigates some risks, it cannot fully eliminate the potential for undesirable outputs. Consequently, post-hoc safety alignment techniques—such as reinforcement learning from human or AI feedback—have become a widely adopted paradigm (dai2023safe; yuan2023gpt; hsu2024safe; yuan-etal-2025-refuse).
However, alignment introduces inherent trade-offs, often conceptualized as the “alignment tax”. Increasing alignment strength can inadvertently constrain the model’s fluency, creativity, or general problem-solving ability. Prior studies have shown that over-aligned models may become verbose, evasive, and less effective at expressive or imaginative tasks (wen2025know; chen2025persona).
Our work extends this line of inquiry to the specific context of persona simulation, empirically validating what we call the “Too Good to be Bad” phenomenon: safety alignment suppresses the model’s capacity to generate convincing portrayals of morally ambiguous or villainous characters. While alignment is beneficial for avoiding real-world harms, it constrains the fidelity of performance in fictional or creative settings where negative traits such as deceit, selfishness, or cruelty are essential to character authenticity. To the best of our knowledge, this study represents the first benchmark-scale evaluation to isolate and quantify this trade-off within the domain of character embodiment, revealing a key limitation of current alignment methodology.
6 Conclusion
In this work, we introduced the Moral RolePlay benchmark to systematically investigate the ability of LLMs to portray characters across the moral spectrum. Our central finding is that state-of-the-art models, while proficient at simulating benevolent figures, exhibit a significant and consistent decline in fidelity when tasked with role-playing antagonistic characters. This failure is not random but is rooted in a conflict with their core safety alignment; models struggle most with traits like deceit, manipulation, and selfishness, which are antithetical to the principles of helpfulness and honesty. We further demonstrated that general conversational prowess is not a reliable indicator of this specialized creative capability.
The implications of our findings extend beyond narrative generation. The inability to simulate the full range of human behaviors, including negative ones, points to a limitation in a model’s understanding of social dynamics and psychology. This highlights a critical trade-off between ensuring safety and achieving high-fidelity representation, which has consequences for applications in the social sciences, education, and art. Future work should focus on developing more sophisticated, context-aware alignment techniques that can distinguish between generating harmful content and simulating fictional antagonism. Our dataset provides a valuable resource for training and evaluating such models, paving the way for LLMs that are both safe and capable of exploring the complete, complex tapestry of human nature.
Appendix A Detailed Results of Robustness Validation
Figure 4: Performance of third-person (default) and first-person roleplay.
Table 11: Impact of reasoning on role-playing quality.
Model Reasoning Level 1 Level 2 Level 3 Level 4
gemini-2.5-pro × 3.39 3.43 2.93 2.80
$\checkmark$ 3.42 3.27 2.97 2.75
claude-opus-4.1 × 3.23 3.08 2.63 2.48
$\checkmark$ 3.20 3.02 2.54 2.53
claude-sonnet-4.5 × 3.20 3.37 2.89 2.56
$\checkmark$ 3.35 3.23 2.83 2.54
qwen3-max × 3.19 2.80 2.53 2.48
$\checkmark$ 2.83 2.83 2.18 2.33
grok-4-fast × 3.15 2.98 2.53 2.43
$\checkmark$ 3.25 3.09 2.67 2.50
deepseek-v3.1 × 3.32 3.16 2.88 2.71
$\checkmark$ 3.37 3.12 2.88 2.82
glm-4.5 × 3.15 3.19 2.81 2.69
$\checkmark$ 3.17 3.10 2.73 2.55